OpenStudy (anonymous):

What are the solutions for x in the proportion 2x+3/x = 2x/2.? A.) x = −1 and x = 3 B.) x = 1 and x = −3 C.) x = −2 and x = 3 D.) x = 2 and x = −3

OpenStudy (mandre):

I think you equation is wrong.

OpenStudy (anonymous):

no its right

OpenStudy (mandre):

*your Can you use the Equation tool?

OpenStudy (anonymous):

\[2x +3divx =2x \div2\]

OpenStudy (mandre):

2x+3/x = 2x/2 multiply by x 2x^2 + 3 = 2x^2 / 2 2x^2 + 3 = x^2 2x^2 + 3 - x^2 = 0 x^2 = -3 x = Sqrt(-3) Can't square root a negative for your purpose. That is if your equation is this: \[2x + \frac{ 3 }{ x } = \frac{ 2x }{ 2 }\]

OpenStudy (anonymous):

so is it d

OpenStudy (mandre):

Multiply by x and simplify \[2x ^{2}+3 = x ^{2}\] \[2x ^{2}+3-x ^{2}=0\] \[x ^{2}=-3\] No. Your equation is wrong.

OpenStudy (mandre):

Take square root both sides and you get Sqrt (-3). Can't get an answer that way.

OpenStudy (anonymous):

ok I c what ur saying thank u now I no what happened